OpenStudy (aroub):

evaluate the limit: lim x-> pi/3 (p(x)+x^2-sqrt3) if lim p(x)=sqrt3 Answer= 1.0966??

OpenStudy (anonymous):

\(\frac{\pi}{9}\)

OpenStudy (rational):

\[\large \lim\limits_{x\to \pi/3} \left(\sqrt{3} + x^2 - \sqrt{3}\right) = \lim\limits_{x\to \pi/3} x^2 = \pi^2/9 \]

OpenStudy (anonymous):

I mean, yes your answer is very correct..

OpenStudy (anonymous):

That is \(\frac{\pi^2}{9}\)

OpenStudy (anonymous):

@aum use \\ to enter. :P

OpenStudy (aum):

\[ \lim_{x \rightarrow \pi/3}(p(x)+x^2-\sqrt3) = \lim_{x \rightarrow \pi/3}p(x)+ \lim_{x \rightarrow \pi/3}x^2- \lim_{x \rightarrow \pi/3}\sqrt3 = \\ \sqrt{3} + \pi^2/9 - \sqrt{3} = \pi^2/9 \]

OpenStudy (anonymous):

That looks great.. :)

OpenStudy (aroub):

Thank youu :D I have few more, if you guys can help? :)

OpenStudy (nincompoop):

but you are suppose to show your solution to us...

OpenStudy (anonymous):

Sure, you write it, we help you.. :)